User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 9th November

Chair: Ross

Present: Jamie, Bryn, Tiege, Patrick, Jesse, Liz, Arwin

Apologies: Jonny, Lucas, Lucia


General Issues:

  1. Calendar didn't include Hobart even though it was on the master schedule. The 26m was a late addition to the AUA schedule. Should be avoided in future.
    1. Let Jamie know if you don't have access rights to the calendar.
    2. nobs looks at master schedule. Warren's script already checks master schedule.
    3. CHECK THE MASTER SCHEDULE! and the start time.
  2. Jesse sent around an email about script. Works for ho, hb, ke and yg, and opens most of the windows needed.
    1. didn't put delay plots in the right position; this is because they were already open.
    2. run windex before windows if there is stuff open.

Handover Issues:

  1. Ho: iread power levels are a bit low 28000 and 38000 of 48000 | changing attentuation did nothing to help (OHG109)
    1. mot an issue: got a different frequency range for O'Higgins (uses Mark 4 rack), slightly smaller range, different power and temperatures. Not paying attention to read in this experiment. Needs more tests.
    2. Tsys temperatures below 200 now seem okay, probably degradation of the system. Measured wrong due to absolute calibration error but maybe 10-20% due to the degradation.
  2. Ho: problems with cryo temp (OHG109 and AUA031)
    1. Just summer maybe. Also receiver pressure a bit dodgy, at 4 microns. Build of gases, got over 5 microns last night but was pumped down. Summer season so cryo problems likely to be more of an issue. Keep an eye out.
  3. Ke: Ran fivept several times, the offsets are looking a bit dodgy (AUA031)
    1. Problem after bad storms in Katherine. Bryn could get below 0.04 offset even after several attempts. Jamie to work on this.
    2. Similarly, fringe testing is quite rubbish these days.
  4. Yg: alarms going off for illegal state error and unable to read systemclock1. antenna=open antenna=operate fixed the issue (R1816)
    1. Bryn breaks stuff. This issue has come up for both Tiege and Liz too. HMI clock get confused about the time; need to reset if the simple fix fails.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_11_09.txt · Last modified: 2017/11/09 02:24 by Ross Turner